# Synthetic-Powered Predictive Inference

### v1

@@ 4.2 Regression on the MEPS dataset
Figure 3 presents the coverage and interval length results for OnlyReal, OnlySynth, and SPPI at target coverage level $0.91-\alpha=$ 0.9, across different age-groups. Similarly to the classification experiments, OnlyReal attains valid coverage but has a higher variance due to the small size of the real calibration set. Following that figure, we see that the synthetic and real data are well aligned. However, OnlySynth, which relies solely on synthetic data, lacks formal coverage guarantees. In contrast, SPPI achieves coverage close to the nominal level of 0.9, as predicted by Theorem 3.3. We provide results for additional $\alpha$ levels and further experiments in Appendix J.
Figure 3: Results for the MEPS dataset: Marginal coverage and interval length for each age-group, obtained by OnlyReal, OnlySynth, and SPPI. Target coverage is $0.91-\alpha=$ 0.9; experiments are repeated for 100 trials.
@@ Appendix A Algorithmic details
Algorithm 1 Synthetic-powered predictive inference (SPPI)
@@
Algorithm 2 SPPI with data-dependent $k$-nearest subset selection
@@ Appendix B Related work
The concept of prediction sets dates back to foundational works such as Wilks [58], Wald [57], Scheffe and Tukey [47], and Tukey [51, 52]. The initial ideas behind conformal prediction were introduced by Saunders et al. [46] and Vovk et al. [54]. Since then, with the rise of machine learning, conformal prediction has emerged as a widely used framework for constructing distribution-free prediction sets [e.g., 37, 56, 53, 11, 15, 31, 30, 32, 33, 45, 19, 38, 4, 23, 21, 22, 43, 9, 16, 34, 35].
@@
Another related line of work is few-shot conformal prediction [18, 39], which addresses settings where only limited data is available for the target task, along with additional auxiliary tasks. These approaches leverage related but distinct tasks to improve sample efficiency. Fisch et al. [18] provides asymptotic task-conditional coverage guarantees, whereas our focus is on finite-sample guarantees. Park et al. [39] mitigate the small-sample challenge using cross-validation, but their methods remain constrained by the overall number of available datapoints—which we assume to be small in our setting.
@@ Appendix D Explaining the score transporter
In this section, we provide further intuition about our proposed procedure, as well as the theoretical bounds established in Theorem 3.5.
A high-level illustration of the proposed method is provided in Figure S1. The illustration outlines the main components and their flow. Given a limited labeled real dataset and a large labeled synthetic dataset, we apply the predictive model to compute prediction errors, referred to as nonconformity scores. We then use the score transporter to map the real scores to the synthetic domain. Using these mapped scores along with the synthetic scores, we perform calibration and construct a prediction set or interval. Next, we provide a more detailed description of the score transporter.
Figure S1: A high-level overview of the proposed method. The approach leverages a small labeled real dataset alongside a large labeled synthetic dataset. The score transporter maps prediction errors (scores) from the real domain to the synthetic one. Calibration is then performed using the transported real errors (scores) and the synthetic errors (scores).
Figure S2 provides a schematic overview of the procedure. For clarity, we illustrate the construction of the prediction set as described in Equation (9). In practice, we use the computationally efficient procedure described in Equation (10), which we have shown to be equivalent in Proposition 3.2.
We begin by computing nonconformity scores for both real and synthetic data points. We assume access to a fixed score function $s$ (e.g., a pre-trained black-box model), along with a small real calibration set ${(X_{i},Y_{i})}_{i\in[m]}$, a large synthetic calibration set ${(\tilde{X}_{j},\tilde{Y}_{j})}_{j\in[N]}$, and a test point $X_{m+1}$.
For each candidate label $y\in\mathcal{Y}$, we compute the nonconformity score of the test pair $(X_{m+1},y)$, along with those of the real and synthetic calibration sets. These scores are depicted as circles in Figure S2, displayed in sorted order. Crucially, the scores for the real and synthetic calibration sets are computed once, and then reused for all candidate labels.
Next, following Equation (7), each real score $S_{(i)}$, $i\in[m+1]$, is associated with a window in the synthetic score space. These windows (illustrated as rectangles connected to each real score) define the set of synthetic scores to which each real score can potentially be mapped. Importantly, these windows depend only on the sample sizes $m$, $N$, and the parameter $\beta$—not on the values of the scores themselves.
For the test score $s(X_{m+1},y)$, we identify its synthetic neighbor after mapping, within its associated window according to (8), denoted by $T(s(X_{m+1},y))$. This synthetic score is then compared to the $(1-\alpha)$th empirical quantile of the synthetic scores, $\tilde{Q}_{1-\alpha}$ (highlighted in blue). If $T(s(X_{m+1},y))\leq\tilde{Q}_{1-\alpha}$, the candidate label $y$ is included in the prediction set $\smash{\widehat{C}(X_{m+1})}$; otherwise, it is excluded.
(a) Candidate label $y_{1}$: the test score (red) ranks second among the real scores. Its synthetic neighbor after mapping (also red) lies below $\tilde{Q}_{1-\alpha}$, so $y_{1}\in\widehat{C}(X_{m+1})$.
(b) Candidate label $y_{2}$: the test score (purple) ranks fourth among the real scores. Its synthetic neighbor after mapping (also purple) lies above $\tilde{Q}_{1-\alpha}$, so $y_{2}\notin\widehat{C}(X_{m+1})$.
Figure S2: Illustration of the synthetic-powered predictive inference for two candidate labels. Each panel shows sorted nonconformity scores: real scores on the left and synthetic scores on the right. The rectangles indicate the windows in the synthetic space to which each real score can be mapped. The blue circle denotes the $(1-\alpha)$th empirical quantile of the synthetic scores, $\tilde{Q}_{1-\alpha}$.
To illustrate this, Figure 2(a) shows the procedure for candidate label $y_{1}\in\mathcal{Y}$. The corresponding test score (marked in red) ranks second among the real scores, and its mapped synthetic neighbor within the corresponding window (also marked in red) lies below the quantile threshold: $T(s(X_{m+1},y_{1}))\leq\tilde{Q}_{1-\alpha}$. Thus, $y_{1}\in\widehat{C}(X_{m+1})$. In contrast, Figure 2(b) depicts the case for label $y_{2}\in\mathcal{Y}$ (marked in purple), where the test score ranks fourth, and its mapped neighbor exceeds the threshold: $T(s(X_{m+1},y_{2}))>\tilde{Q}_{1-\alpha}$. Therefore, $y_{2}\notin\widehat{C}(X_{m+1})$. This procedure is repeated for each $y\in\mathcal{Y}$ to construct the full prediction set $\smash{\widehat{C}(X_{m+1})}$.
We now illustrate the theoretical bounds established in Theorem 3.5, using the same schematic from Figure S2. Real and synthetic nonconformity scores are shown as circles in sorted order, with each real score connected to a window in the synthetic score space (depicted as rectangles). The $(1-\alpha)$th empirical quantile of the synthetic scores, $\tilde{Q}_{1-\alpha}$, is marked by a blue circle.
@@
(a) Values of $R^{+}_{r}$ (in red), representing the upper endpoints of the synthetic windows. The smallest three real scores satisfy $R^{+}_{r}\leq\tilde{Q}_{1-\alpha}$ and are thus guaranteed to be mapped to synthetic scores below the threshold $\tilde{Q}_{1-\alpha}$.
(b) Values of $R^{-}_{r}$ (in green), representing the lower endpoints of the synthetic windows. The fifth real score satisfies $R^{-}_{(5)}>\tilde{Q}_{1-\alpha}$, meaning it is necessarily mapped above the threshold $\tilde{Q}_{1-\alpha}$.
@@
Figure 3(a) shows the values $R^{+}_{r}$ (in red), which are used to compute the lower bound. Since the transported score $T(S_{(r)})$—defined as the nearest synthetic score within the window among those that are smaller than $S_{(r)}$—is always less than or equal to $R^{+}_{r}$, any real score satisfying $R^{+}_{r}\leq\tilde{Q}_{1-\alpha}$ must necessarily satisfy $T(S_{(r)})\leq\tilde{Q}_{1-\alpha}$. Consequently, the corresponding label will always be included in the prediction set.
Figure 3(b) shows the values $R^{-}_{r}$ (in green), which are used to compute the upper bound. Real scores for which $R^{-}_{r}\leq\tilde{Q}_{1-\alpha}$ may be mapped to a synthetic score below the threshold and thus may be included in the prediction set—for example, the bottom four real scores in the figure. In contrast, if $R^{-}_{r}>\tilde{Q}_{1-\alpha}$ (as for the fifth score), then the transported score must exceed the threshold, and the corresponding label is guaranteed to be excluded.
@@ Appendix F Predictive inference with label-conditional coverage control
Here, we review the standard approach [56] for achieving the label-conditional coverage guarantee (3), and then discuss a variant of this approach based on SPPI.
@@
Now, we introduce the SPPI-based method that ensures the label-conditional coverage guarantee. The idea follows the same logic as the standard method: We run SPPI within each class-specific partition of the real and synthetic calibration sets and then combine the results. For a class that does not appear in the synthetic dataset, we run the procedure with the entire synthetic dataset.
@@ I.1 Experiments with generated synthetic data
Figure S8 presents the performance under both marginal and label-conditional guarantees at levels $\alpha=0.02$ and $0.1$. We observe a similar trend to that seen in Figure 1. Following that figure, we can see that the standard conformal prediction method, OnlyReal, controls the coverage at the $1-\alpha$ level as expected, but it produces overly conservative prediction sets. OnlySynth method fails to achieve the target coverage level of $1-\alpha$, under-covering some classes, while in others, it becomes overly conservative, depending on the unknown distribution shift between the real and synthetic data. In contrast, the proposed method, SPPI, stays within the theoretical bounds and produces informative prediction sets.
@@
Figure S8: Results for the ImageNet data: Coverage rates of OnlyReal, OnlySynth, and SPPI run at level $\alpha=0.02$ (a) and $0.1$ (b), averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Tables S2 and S4 for results across all classes.
Tables S3, S2 and S4 present results for all 30 classes in the real calibration set corresponding to Figure 1 in the main manuscript and Figures 8(a) and 8(b) above.
Table S2: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.981-\alpha=$ 0.98. The theoretical coverage guarantees for SPPI are in the range $[93.7,100]$. Other details are as in Figure S8.
@@
Class Only Real Only Synth SPPI Only Real Only Synth SPPI
@@
Table S3: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.951-\alpha=$ 0.95. The theoretical coverage guarantees for SPPI are in the range $[93.7,100]$. Other details are as in Figure 1.
@@
Class Only Real Only Synth SPPI Only Real Only Synth SPPI
@@
Table S4: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.91-\alpha=$ 0.9. The theoretical coverage guarantees for SPPI-Subset are in the range $[81.2,93.7]$. Other details as in Figure S8.
@@
Class Only Real Only Synth SPPI Only Real Only Synth SPPI
@@ I.1.1 The effect of the real calibration set size
Here, we evaluate the performance of different methods as a function of the real calibration set size $m$, following the same setup described in Section 4.1.1. This parameter directly affects the performance of both the standard conformal prediction method, OnlyReal, and our proposed method, SPPI, including the theoretical bounds established in Theorem 3.5. In contrast, OnlySynth, which relies solely on the synthetic calibration set, is unaffected by changes in $m$. As such, it serves as a useful baseline for assessing how well the synthetic calibration set aligns with the real one.
@@
Figure S9: Results for the ImageNet data: Coverage rate for OnlyReal, OnlySynth, and SPPI on the “Lighter” class as a function of the real calibration set size $m$, for levels $\alpha=0.02$ (a), $\alpha=0.05$ (b), and $\alpha=0.1$ (c).
@@
In contrast, our proposed method, SPPI, achieves coverage within the theoretical bounds even for small $m$, with reduced variance in coverage and smaller prediction sets for $m\geq 15$. Interestingly, for $\alpha=0.02$ and $m=5$ or $10$, the theoretical lower and upper coverage bounds are both equal to unity, indicating that we know a priori that the proposed method yields trivial prediction sets for this window construction.
For $\alpha=0.05$ and $\alpha=0.1$ (Figures 9(b) and 9(c), respectively), we observe similar trends. Our method, SPPI, consistently achieves coverage within the theoretical bounds, remaining close to the target coverage level $1-\alpha$, while also exhibiting reduced variance in coverage and producing smaller, more informative prediction sets compared to the baseline, OnlyReal.
Additionally, Figure S10 presents the same experiment as in Figure S9, but for the “Beaver” class. In this case, the OnlySynth method yields coverage that falls significantly below the target level $1-\alpha$, indicating that the synthetic calibration set differs substantially from the real data. Nevertheless, our proposed method, SPPI, achieves coverage within the theoretical bounds across all $\alpha$ levels and calibration set sizes, while also producing informative prediction sets.
@@
Figure S10: Results for the ImageNet data: Coverage rate for OnlyReal, OnlySynth, and SPPI on the Beaver class as a function of the real calibration set size $m$, for levels $\alpha=0.02$ (a), $\alpha=0.05$ (b), and $\alpha=0.1$ (c).
I.1.2 Results for SPPI with FLUX-generated synthetic data
In this section, we evaluate the performance of the proposed method, SPPI, using synthetic images generated by the FLUX.1 model [28]. The experimental setup follows the same procedure described in Section 4.1.1 of the main manuscript. As before, we aim for both marginal and label-conditional coverage guarantees.
Figure S11 presents the marginal and label-conditional coverage of various methods at levels $\alpha=0.02$, $0.05$, and $0.1$. The results for label-conditional guarantees are presented for representative classes; results for all classes in the real population are detailed in Tables S5, S6 and S7. We observe similar trends to those observed using synthetic images generated by Stable Diffusion. The standard conformal method, OnlyReal, controls the coverage at the $1-\alpha$ level; however, it yields overly conservative prediction sets due to the small sample size. OnlySynth fails to control the coverage at the desired level, exhibiting under-coverage for some classes and over-coverage for others. In contrast, the proposed method, SPPI, achieves coverage within the theoretical bounds while providing smaller, more informative prediction sets.

### v2

@@ 4.2 Regression on the MEPS dataset
Figure 6 presents the coverage and interval length results for OnlyReal, OnlySynth, and SPI at target coverage level $0.91-\alpha=$ 0.9, across different age-groups. Similarly to the classification experiments, OnlyReal attains valid coverage but has a higher variance due to the small size of the real calibration set. Following that figure, we see that the synthetic and real data are well aligned. However, OnlySynth, which relies solely on synthetic data, lacks formal coverage guarantees. In contrast, SPI achieves coverage close to the nominal level of 0.9, as predicted by Theorem 3.3. We provide results for additional $\alpha$ levels and further experiments in Appendix J.
Figure 6: Results for the MEPS dataset: Marginal coverage and interval length for each age-group, obtained by OnlyReal, OnlySynth, and SPI. Target coverage is $0.91-\alpha=$ 0.9; experiments are repeated for 100 trials.
@@ Appendix A Algorithmic details
Algorithm 1 Synthetic-powered predictive inference (SPI)
@@
Algorithm 2 SPI with data-dependent $k$-nearest subset selection
@@ Appendix B Related work
The concept of prediction sets dates back to foundational works such as Wilks [60], Wald [59], Scheffe and Tukey [48], and Tukey [53, 54]. The initial ideas behind conformal prediction were introduced by Saunders et al. [47] and Vovk et al. [56]. Since then, with the rise of machine learning, conformal prediction has emerged as a widely used framework for constructing distribution-free prediction sets [e.g., 38, 58, 55, 11, 15, 32, 31, 33, 34, 46, 20, 39, 4, 24, 22, 23, 44, 9, 17, 35, 36].
@@
Another related line of work is few-shot conformal prediction [19, 40], which addresses settings where only limited data is available for the target task, along with additional auxiliary tasks. These approaches leverage related but distinct tasks to improve sample efficiency. Fisch et al. [19] provides asymptotic task-conditional coverage guarantees, whereas our focus is on finite-sample guarantees. Park et al. [40] mitigate the small-sample challenge using cross-validation, but their methods remain constrained by the overall number of available datapoints—which we assume to be small in our setting. Dutta et al. [16] propose to retrieve web images to enable conformal prediction in zero-shot settings, by leveraging conformal prediction with ambiguous ground truth [51], but do not provide coverage guarantees for their method.
@@ Appendix D Explaining the score transporter
In this section, we provide further intuition about the theoretical bounds established in Theorem 3.5.
We illustrate these bounds using the same schematic from Figure 2. Real and synthetic nonconformity scores are shown as circles in sorted order, with each real score connected to a window in the synthetic score space (depicted as rectangles). The $(1-\alpha)$th empirical quantile of the synthetic scores, $\tilde{Q}_{1-\alpha}$, is outlined in black.
@@
(a) Values of $R^{+}_{r}$ (in blue), representing the upper endpoints of the synthetic windows. The smallest three real scores satisfy $R^{+}_{r}\leq\tilde{Q}_{1-\alpha}$ and are thus guaranteed to be mapped to synthetic scores below the threshold $\tilde{Q}_{1-\alpha}$.
(b) Values of $R^{-}_{r}$ (in blue), representing the lower endpoints of the synthetic windows. The fifth-ranked real score satisfies $R^{-}_{5}>\tilde{Q}_{1-\alpha}$, meaning it is necessarily mapped above the threshold $\tilde{Q}_{1-\alpha}$.
@@
Figure 1(a) shows the values $R^{+}_{r}$ (in blue), which are used to compute the lower bound. Since the transported score $T(S_{(r)})$—defined as the nearest synthetic score within the window among those that are smaller than $S_{(r)}$—is always less than or equal to $R^{+}_{r}$, any real score satisfying $R^{+}_{r}\leq\tilde{Q}_{1-\alpha}$ must necessarily satisfy $T(S_{(r)})\leq\tilde{Q}_{1-\alpha}$. Consequently, the corresponding label will always be included in the prediction set.
Figure 1(b) shows the values $R^{-}_{r}$ (in blue), which are used to compute the upper bound. Real scores for which $R^{-}_{r}\leq\tilde{Q}_{1-\alpha}$ may be mapped to a synthetic score below the threshold and thus may be included in the prediction set—for example, the bottom four real scores in the figure. In contrast, if $R^{-}_{r}>\tilde{Q}_{1-\alpha}$ (as for the fifth-ranked score), then the transported score must exceed the threshold, and the corresponding label is guaranteed to be excluded.
@@ Appendix F Predictive inference with label-conditional coverage control
Here, we review the standard approach [58] for achieving the label-conditional coverage guarantee (3), and then discuss a variant of this approach based on SPI.
@@
Now, we introduce the SPI-based method that ensures the label-conditional coverage guarantee. The idea follows the same logic as the standard method: We run SPI within each class-specific partition of the real and synthetic calibration sets and then combine the results. For a class that does not appear in the synthetic dataset, we run the procedure with the entire synthetic dataset.
@@ I.1 Experiments with generated synthetic data
Figure S6 presents the performance under both marginal and label-conditional guarantees at levels $\alpha=0.02$ and $0.1$. We observe a similar trend to that seen in Figure 4. Following that figure, we can see that the standard conformal prediction method, OnlyReal, controls the coverage at the $1-\alpha$ level as expected, but it produces overly conservative prediction sets. OnlySynth method fails to achieve the target coverage level of $1-\alpha$, under-covering some classes, while in others, it becomes overly conservative, depending on the unknown distribution shift between the real and synthetic data. In contrast, the proposed method, SPI, stays within the theoretical bounds and produces informative prediction sets.
@@
Figure S6: Results for the ImageNet data: Coverage rates of OnlyReal, OnlySynth, and SPI run at level $\alpha=0.02$ (a) and $0.1$ (b), averaged over 100 trials. Left: Average coverage. Right: Average prediction set size, both under marginal (leftmost box in each group) and label-conditional coverage settings. Label-conditional results are shown for selected classes; see Tables S2 and S4 for results across all classes.
Tables S3, S2 and S4 present results for all 30 classes in the real calibration set corresponding to Figure 4 in the main manuscript and Figures 6(a) and 6(b) above.
Table S2: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.981-\alpha=$ 0.98. The theoretical coverage guarantees for SPI are in the range $[93.7,100]$. Other details are as in Figure S6.
@@
Class Only Real Only Synth SPI Only Real Only Synth SPI
@@
Table S3: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.951-\alpha=$ 0.95. The theoretical coverage guarantees for SPI are in the range $[93.7,100]$. Other details are as in Figure 4.
@@
Class Only Real Only Synth SPI Only Real Only Synth SPI
@@
Table S4: Per-class conditional coverage (in %) and prediction set size for each method, computed over 100 trials. Standard errors are shown in parentheses. The target coverage level is $0.91-\alpha=$ 0.9. The theoretical coverage guarantees for SPI are in the range $[81.2,93.7]$. Other details as in Figure S6.
@@
Class Only Real Only Synth SPI Only Real Only Synth SPI
@@ I.1.1 The effect of the real calibration set size
Here, we evaluate the performance of different methods as a function of the real calibration set size $m$, following the same setup described in Section 4.1.1. This parameter directly affects the performance of both the standard conformal prediction method, OnlyReal, and our proposed method, SPI, including the theoretical bounds established in Theorem 3.5. In contrast, OnlySynth, which relies solely on the synthetic calibration set, is unaffected by changes in $m$. As such, it serves as a useful baseline for assessing how well the synthetic calibration set aligns with the real one.
@@
Figure S7: Results for the ImageNet data: Coverage rate for OnlyReal, OnlySynth, and SPI on the “Lighter” class as a function of the real calibration set size $m$, for levels $\alpha=0.02$ (a), $\alpha=0.05$ (b), and $\alpha=0.1$ (c).
@@
In contrast, our proposed method, SPI, achieves coverage within the theoretical bounds even for small $m$, with reduced variance in coverage and smaller prediction sets for $m\geq 15$. Interestingly, for $\alpha=0.02$ and $m=5$ or $10$, the theoretical lower and upper coverage bounds are both equal to unity, indicating that we know a priori that the proposed method yields trivial prediction sets for this window construction.
For $\alpha=0.05$ and $\alpha=0.1$ (Figures 7(b) and 7(c), respectively), we observe similar trends. Our method, SPI, consistently achieves coverage within the theoretical bounds, remaining close to the target coverage level $1-\alpha$, while also exhibiting reduced variance in coverage and producing smaller, more informative prediction sets compared to the baseline, OnlyReal.
Additionally, Figure S8 presents the same experiment as in Figure S7, but for the “Beaver” class. In this case, the OnlySynth method yields coverage that falls significantly below the target level $1-\alpha$, indicating that the synthetic calibration set differs substantially from the real data. Nevertheless, our proposed method, SPI, achieves coverage within the theoretical bounds across all $\alpha$ levels and calibration set sizes, while also producing informative prediction sets.
@@
Figure S8: Results for the ImageNet data: Coverage rate for OnlyReal, OnlySynth, and SPI on the Beaver class as a function of the real calibration set size $m$, for levels $\alpha=0.02$ (a), $\alpha=0.05$ (b), and $\alpha=0.1$ (c).
I.1.2 Results for SPI with FLUX-generated synthetic data
In this section, we evaluate the performance of the proposed method, SPI, using synthetic images generated by the FLUX.1 model [29]. The experimental setup follows the same procedure described in Section 4.1.1 of the main manuscript. As before, we aim for both marginal and label-conditional coverage guarantees.
Figure S9 presents the marginal and label-conditional coverage of various methods at levels $\alpha=0.02$, $0.05$, and $0.1$. The results for label-conditional guarantees are presented for representative classes; results for all classes in the real population are detailed in Tables S5, S6 and S7. We observe similar trends to those observed using synthetic images generated by Stable Diffusion. The standard conformal method, OnlyReal, controls the coverage at the $1-\alpha$ level; however, it yields overly conservative prediction sets due to the small sample size. OnlySynth fails to control the coverage at the desired level, exhibiting under-coverage for some classes and over-coverage for others. In contrast, the proposed method, SPI, achieves coverage within the theoretical bounds while providing smaller, more informative prediction sets.